# New Insights Into Achievable Spectral Efficiency With Adaptive Free Space Optical Transmissions

### v8

@@ Abstract
Terrestrial free-space optical (FSO) communication systems, while designed to operate on large unlicensed optical bandwidths, are power-constrained due to strict eye safety regulations. The channel fluctuation inherent in terrestrial FSO links also limits the received optical power. Consequently, the available signal-to-noise ratio (SNR) per Hz could become limited; this holds for future long-haul terrestrial systems based on coherent optical communications. An efficient and adaptive transmission mechanism is thus crucial at the optical source. However, a critical assessment of the impact of adaptive transmission in terrestrial FSO systems has received less attention in the literature. This work studies terrestrial FSO communication systems employing adaptive beam transmission while detection receivers operate under shot noise-limited conditions. Specifically, we propose a novel exact closed-form expression for the average spectral efficiency of these FSO systems with transmit beam power adaptation over the gamma-gamma turbulence channel with pointing error. More importantly, we provide a detailed assessment of the impact of turbulence and pointing error impairments on the performance of the aforementioned FSO systems, revealing several novel and counterintuitive insights. In particular, the extensive numerical results help elucidate the intricacies of analyzing these terrestrial FSO systems and clarify a few misconceptions alluded to in recent related literature.
Index Terms:
Average spectral efficiency, free-space optical (FSO) communications, optical beam power adaptation, gamma-gamma turbulence, pointer error, shot noise.
@@ I Introduction
(ii) Coherent heterodyne detection (HD): Coherent systems exploit both the amplitude and phase of the optical carrier to convey information, enabling two degrees of freedom. In heterodyne detection, a locally generated optical signal—typically offset in frequency from the received signal—is mixed with the incoming beam prior to photodetection. This approach significantly improves receiver sensitivity, which can be leveraged to extend link distances. Additionally, coherent receivers enable high selectivity via narrowband RF filtering, in contrast to the broader optical filtering in IM-DD systems [10].
@@
Transmit optical power in terrestrial FSO communications is limited by stringent eye and skin safety regulations [20]. In contrast, FSO links benefit from virtually unlimited optical bandwidths (tens to hundreds of gigahertz per wavelength) to support ultra-high-speed communications [12], making the available power per hertz a potential limiting factor. While such low-SNR operating points are rarely encountered today, recent demonstrations of extreme-wideband wavelength-division multiplexing (WDM) coherent FSO systems indicate that distributing a fixed transmit power over many carriers may plausibly lead to lower per-channel SNR [7], [21]. In addition, variability in atmospheric attenuation under haze, fog, or pollution can accentuate this effect. Thus, operation in the low-SNR regime may emerge as a relevant consideration for future ultra-wideband WDM terrestrial FSO systems. In addition, irradiance fluctuations (mentioned earlier) present in terrestrial FSO channels degrade the power collected at the receiver [11], but this degradation can be alleviated through power control. For instance, a simple scheme in which the transmitter activates only when the instantaneous channel quality exceeds a predefined threshold can be effective under low-power budgets. Importantly, typical terrestrial FSO channels exhibit extremely high coherence bandwidths (hundreds of gigahertz or more) [22, 23] and relatively long coherence times on the order of milliseconds [24]. These conditions yield a quasi-static, frequency-flat block-fading channel, which is increasingly exploited by recent field trials [6, 7, 8, 21] for adaptive transmission in terrestrial coherent FSO links. This context motivates the subsequent analysis of spectral efficiency in low-SNR regimes, where turbulence-induced fluctuations and limited per-Hz transmit power, especially in ultra-wideband WDM coherent FSO systems, can meaningfully affect achievable throughput.
@@
In contrast to the recent practical advances mentioned above, adaptive terrestrial FSO communication systems have attracted comparatively limited rigorous theoretical analysis in the literature, as summarized next. Prior works on IM-DD and coherent HD based terrestrial FSO communication systems have examined achievable spectral efficiency (SE) under log-normal and gamma-gamma (GG) turbulence models (e.g., [25, 26, 27, 28, 29, 30], and references therein). More recent studies have extended this analysis to generalized turbulence channels, such as those modeled by the double-generalized gamma and Málaga ($\mathcal{M}$) distributions [31], [32], [33], [34]. Despite these contributions, the critical aspect of beam power adaptation at the optical transmitter has been largely overlooked-except in a few cases, such as [25], [26], and [27]. This omission is likely due to the analytical complexity involved in solving the associated optimization problems. In [26], an infinite series solution for the SE of an HD-based adaptive FSO system operating over GG turbulence is presented; however, this approach introduces concerns related to truncation error and interpretability. Exact and asymptotic SE expressions for IM-DD-based adaptive systems under thermal noise-limited GG channels are derived in [25][27]. Notably, these works do not address the inherent non-convexity of the SE maximization problem in this regime (see Remark 2), a critical oversight that affects the validity of the reported results. Furthermore, the authors in [32] present a unified SE expression for IM-DD and HD based FSO communication under the $\mathcal{M}$ distribution, which includes the GG distribution as a special case; however, this analysis assumes a fixed transmit beam power, without adaptation. It is also important to note that several FSO related studies [25], [28], [29], [31], [32] lack clarity in the methodology used to simulate varying turbulence conditions. In terrestrial FSO systems, several channel parameters are interdependent-for example, changes in the FSO link length simultaneously affect turbulence severity level, path loss, and pointing error statistics. A rigorous understanding of these interrelations grounded in physical modeling is crucial for accurately evaluating FSO system performance. Neglecting such dependencies may lead to misleading interpretations and flawed insights.
@@
To address these research gaps in terrestrial FSO research, we will utilize adaptive optical signal transmission on existing terrestrial FSO communication systems with an emphasis on spectral efficiency improvement, which basically relies on exploiting the temporal channel fluctuations due to turbulence and pointing error impairments. Specifically, this work aims to provide in-depth insights into the operation of terrestrial FSO communication systems, and provide a baseline SE performance for future adaptive FSO systems with strong practical constraints.
The principal contributions of this paper are summarized as follows.
@@
  • We present a novel, unified, and exact closed-form expression for the average spectral efficiency (ASE) of beam power adaptation-based FSO communication systems with IM-DD and HD receivers, accounting for gamma-gamma (GG) turbulence and pointing error impairments. Theorem 1, along with its specialization in Corollary 1 for the case without pointing errors, represents the central analytical contribution of this work.
@@
The remainder of this paper is organized as follows. The optical turbulence and pointing error models adopted for terrestrial FSO channel are presented in Subsection II-A. The relevant noise sources and the considered FSO receiver models are described in Subsections II-B and II-C, respectively. Theorems on the spectral efficiency of the turbulent optical channel are derived in Section III and their extensive numerical analysis is presented in Section IV. Finally, Section V summarizes the main results and provides concluding remarks along with directions for future work.
@@ II Channel and System Models
II-A Gamma–Gamma Turbulence with Pointing Error
@@
where $a$ and $b$ are the shape parameters [14], $\Gamma(\cdot)$ is the Gamma function [36, Eq. (8.310.1)], and $K_{v}(\cdot)$ is the modified Bessel function of the second kind and order $v$ [36, Eq. (8.494.1)]. For plane wave optical radiation at the receiver, the GG parameters $a$ and $b$ are related to atmospheric conditions as
$a\,$ $=\,\Bigl{[}{\exp\bigl{(}{{{0.49\sigma_{\mathrm{R}}^{2}}}/{{{{\bigl{(}{1+1.11\sigma_{\mathrm{R}}^{12/5}}\bigr{)}}^{7/6}}}}}\bigl{)}-1}\Bigr{]}^{-1},$ (3)
$b\,$ $=\,\Bigl{[}{\exp({{{0.51\sigma_{\mathrm{R}}^{2}}}/{{{{\bigl{(}{1+0.69\sigma_{\mathrm{R}}^{12/5}}\bigr{)}}^{5/6}}}}}\bigl{)}-1}\Bigr{]}^{-1},$ (4)
@@
where $\sigma_{\mathrm{R}}^{2}=1.23\,C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ is the Rytov variance determining the atmospheric turbulence strength, $k_{\mathrm{w}}=2\pi/{\lambda_{\mathrm{w}}}$ is the optical wave number, $\lambda_{\mathrm{w}}$} is the wavelength, $L$ is the propagation distance, and $C_{\mathrm{n}}^{2}$ is the index of refraction structure parameter [14].
@@
Pointing Error: Pointing error refers to the misalignment of the laser beam with the optical receiver, caused by i) boresight, the fixed displacement between the beam centroid and the receiver center, and ii) jitter, the random temporal displacement due to factors such as building sway, wind, or vibrations. While careful deployment or fast-tracking transmitters can minimize boresight misalignment, residual jitter remains a concern even with a well-designed ATP subsystem [11]. To account for this effect, the severity of jitter in this work is later categorized into representative regimes reflecting different levels of ATP capability (see Table IV in Section IV).
@@
After propagating a distance $L$, the laser beam with waist $w_{\mathrm{L}}$ is incident on the photodetector with aperture radius $r_{\mathrm{A}}$ [37]. The received optical beam attenuation due to turbulence-induced spread and pointing error, described by radial displacement $r$ between the beam footprint and the detector center, is given by $I_{\mathrm{p}}(r;L)\approx A_{0}\exp(-2r^{2}/w_{\mathrm{L_{eq}}}^{2})$ where $w_{\mathrm{L_{eq}}}^{2}=w_{\mathrm{L}}^{2}\sqrt{\pi}\,\text{erf}(v)/(2v\exp(-v^{2}))$ is the (squared) equivalent beam width with $v=\sqrt{\pi}r_{\mathrm{A}}/\sqrt{2}w_{\mathrm{L}}$. The fraction of the collected power at $r=0$ is given by
$A_{0}=[\mathrm{erf}(v)]^{2},$ (5)
@@
where $\text{erf}(\cdot)$ is the standard error function [36, Eq. (8.250.1)], and $0\leqslant A_{0}\leqslant 1$ [15]. For a coherent beam, the beam radius $w_{\mathrm{L}}$ is related to the beam waist $w_{0}$ at the optical transmitter’s exit aperture by $w_{\mathrm{L}}=w_{0}(1+\epsilon(\lambda_{\mathrm{w}}L/\pi w_{0}^{2})^{2})^{1/2}$ where $\epsilon=(1+2w_{0}^{2}/\rho_{0}^{2}(L))$ is the global coherence parameter, $\rho_{0}(L)=(1.46C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{2}L)^{-3/5}$ is the coherence length in turbulence [38], and the transmit beam waist $w_{0}$ is related to the transmitter aperture diameter $D$ as $w_{0}={D}/{\sqrt{2}\pi}$ [39]. Farid and Hranilovic [15] derived the PDF of the irradiance component $I_{\mathrm{p}}$ under the assumptions of i.i.d. zero-mean Gaussian distributions for both horizontal and vertical receiver sways (i.e., zero boresight and identical jitters), which is expressed as
$\phantom{xxxxxxxxxx}f_{I_{\mathrm{p}}}(I_{\mathrm{p}})\,=\,\tfrac{\xi^{2}}{A_{0}^{\xi^{2}}}I_{\mathrm{p}}^{\xi^{2}-1};\ \ \ 0\leqslant I_{\mathrm{p}}\leqslant A_{0}$ (6)
@@
where the parameter $\xi\triangleq{w_{\mathrm{L_{eq}}}}/{2\sigma_{\mathrm{e}}}$ is the ratio of the received equivalent beam waist $w_{\mathrm{L_{eq}}}$ and the standard deviation of the pointing error displacement (jitter). This model is valid when $w_{L}/r_{A}\gg 1$ and further that the effects of $\xi$ and $A_{0}$ parameters are independent [15].
@@
A recent refinement of the pointing error model is introduced in [40], where multiple alternative parameterizations of $\xi$ and $A_{0}$ are proposed to enhance modeling accuracy and computational efficiency (see [40, Fig. 5]). In this work, we hence borrow the Modified Intensity Uniform Model from [40], which retains the same mathematical form for the probability distribution as in (6), but with redefined parameters given by $A_{0}=1-\exp(-2r_{A}^{2}/w_{L}^{2})$ and $\xi^{2}=r_{A}^{2}/(2\sigma_{\mathrm{e}}^{2}A_{0})$. Precisely, this choice demonstrates more than an order-of-magnitude reduction in normalized mean square error (NMSE) compared to the Farid–Hranilovic formulation (see [40, Table 1]), thereby offering a substantially more reliable characterization of pointing jitter effects in terrestrial FSO links.
@@
$f_{I}(I)=\frac{ab\xi^{2}}{A_{0}\Gamma(a)\Gamma(b)}G_{1,3}^{3,0}\left(\frac{abI}{A_{0}}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,\,a-1,\,b-1\\ \end{array}\hskip-5.0pt\right).$ (7)
@@
II-B Noise Sources in FSO Wireless Systems

### v9

# On Achievable Spectral Efficiency Using Adaptive Transmission Over Terrestrial Coherent FSO Links

@@ Abstract
Terrestrial free-space optical (FSO) communication systems, while designed to operate on large unlicensed optical bandwidths, are power-constrained due to strict eye safety regulations. The channel fluctuation inherent in terrestrial FSO links also limits the received optical power. Consequently, the available signal-to-noise ratio (SNR) per Hz could become limited; this holds for future terrestrial systems based on coherent optical communications. An efficient and adaptive transmission mechanism is thus crucial at the optical transmitter. However, a critical assessment of the impact of adaptive transmission in terrestrial FSO systems has received less attention in the literature. This work studies terrestrial coherent FSO communication systems employing adaptive beam transmission while detection receiver operate under shot noise-limited conditions. Specifically, we propose a novel exact closed-form expression for the average spectral efficiency of a coherent FSO system with optimal adaptive transmissions over the gamma-gamma turbulence channel with pointing errors. More importantly, we provide a detailed assessment of the impact of turbulence and pointing error impairments on the coherent FSO system performance, revealing several novel and counterintuitive insights. In particular, the extensive numerical results help elucidate the intricacies of analyzing these terrestrial FSO systems and clarify a few misconceptions alluded to in recent related literature.
@@ I Introduction
(ii) Coherent heterodyne detection (HD): Coherent FSO communication systems exploit both the amplitude and phase of the optical carrier to convey information, enabling two degrees of freedom. In HD, a locally generated optical signal—typically offset in frequency from the received signal—is mixed with the incoming beam prior to photodetection. This detection approach significantly improves receiver sensitivity. Additionally, coherent receivers enable high selectivity via narrowband RF filtering, in contrast to the broader optical filtering in IM-DD based FSO communication systems [10].
@@
Transmit optical power in terrestrial FSO communications is limited by stringent eye and skin safety regulations [20]. In contrast, FSO links benefit from virtually unlimited optical bandwidths (tens to hundreds of gigahertz per wavelength) to support ultra-high-speed communications [12], making the available power per hertz a potential limiting factor. While such low-SNR operating points are rarely encountered today, recent demonstrations of extreme-wideband wavelength-division multiplexing (WDM) coherent FSO systems indicate that distributing a fixed transmit power over many carriers may plausibly lead to lower per-channel SNR [7], [21]. In particular, variability in atmospheric attenuation under haze, fog, or pollution can accentuate this effect. Thus, operation in the low-SNR regime may emerge as a relevant consideration for future ultra-wideband WDM terrestrial FSO systems. In addition, irradiance fluctuations (mentioned earlier) present in terrestrial FSO channels degrade the power collected at the receiver [11], but this degradation can be alleviated through power control. For instance, a simple scheme in which the transmitter activates only when the instantaneous channel quality exceeds a predefined threshold can be effective under low-power budgets. Importantly, typical terrestrial FSO channels exhibit extremely high coherence bandwidths (hundreds of gigahertz or more) [22, 23] and relatively long coherence times on the order of milliseconds [24]. These conditions yield a quasi-static, frequency-flat block-fading channel, which is increasingly exploited by recent field trials [6, 7, 8, 21] for adaptive transmission in terrestrial coherent FSO links. This context motivates the subsequent analysis of spectral efficiency in low-SNR regimes, where turbulence-induced fluctuations and limited per-Hz transmit power, especially in ultra-wideband WDM coherent FSO systems, can meaningfully affect achievable throughput.
@@
In contrast to the recent practical advances mentioned above, adaptive terrestrial FSO communication systems have attracted comparatively limited rigorous theoretical analysis in the literature, as summarized next. Prior works on IM-DD and coherent HD based terrestrial FSO communication systems have examined achievable spectral efficiency (SE) under log-normal and gamma-gamma (GG) turbulence models (e.g., [25, 26, 27, 28, 29, 30], and references therein). More recent studies have extended this analysis to generalized turbulence channels, such as those modeled by the double-generalized gamma and Málaga ($\mathcal{M}$) distributions [31], [32], [33], [34]. Despite these contributions, the critical aspect of beam power adaptation at the optical transmitter has been largely overlooked—except in a few cases, such as [25], [26], and [27]. This omission is likely due to the analytical complexity involved in solving the associated optimization problems. In a closely related work [26],the SE of an HD-based adaptive FSO system over GG turbulence is examined. However, this analysis does not account for the critical influence of pointing errors, which are particularly significant in terrestrial coherent optical communication systems. Furthermore, [32] presents a unified SE expression for IM-DD and HD based FSO communication systems operating under the $\mathcal{M}$ distribution, which subsumes the GG distribution as a special case. However, this analysis assumes a fixed transmit beam power, without adaptation. It is also important to note that several FSO related studies [25], [28], [29], [31], [32] lack clarity in the methodology used to simulate varying turbulence conditions. In terrestrial FSO systems, several channel parameters are interdependent—for example, changes in the FSO link length simultaneously affect turbulence severity level, path loss, and pointing error statistics. A rigorous understanding of these interrelations grounded in physical modeling is crucial for accurately evaluating FSO system performance. Neglecting such dependencies may lead to misleading interpretations and flawed insights.
@@
To address these research gaps in terrestrial FSO research, we will utilize adaptive optical signal transmission on existing terrestrial FSO communication systems with an emphasis on spectral efficiency improvement, which basically relies on exploiting the temporal channel fluctuations due to turbulence and pointing error impairments. Specifically, this work aims to provide in-depth insights into the operation of terrestrial FSO communication systems, and provide a baseline SE performance for future adaptive coherent FSO systems with strong practical constraints. The principal contributions of this paper are as follows:
@@
  • We present a novel and exact closed-form expression for the average spectral efficiency (ASE) of beam power adaptation-based coherent FSO communication systems with HD receivers, accounting for gamma-gamma (GG) turbulence and pointing error impairments. Theorem 1, along with its specialization in Corollary 1 for the case without pointing errors, represents the central analytical contribution of this work.
@@
The remainder of this paper is organized as follows. The optical turbulence and pointing error models adopted for terrestrial FSO channel are presented in Subsection II-A. The relevant noise source and the coherent FSO system model are described in Subsections II-B and II-C, respectively. Theorems on the spectral efficiency of the turbulent optical channel are derived in Section III and their extensive numerical analysis is presented in Section IV. Finally, Section V summarizes the main results and provides concluding remarks along with directions for future work.
@@ II Channel and System Models
II-A Gamma–Gamma Turbulence with Pointing Errors
@@
where $a$ and $b$ are shape parameters [14], $\Gamma(\cdot)$ is the Gamma function [36, Eq. (8.310.1)], and $K_{v}(\cdot)$ is the modified Bessel function of the second kind and order $v$ [36, Eq. (8.494.1)]. For plane-wave optical radiation at the receiver, the GG shaping parameters depend on atmospheric conditions as
$a\,$ $=\,\bigl[{\exp\bigl({{0.49\sigma_{\mathrm{\scriptscriptstyle R}}^{2}}}{\bigl({1+1.11\sigma_{\mathrm{\scriptscriptstyle R}}^{2.4}}\bigr)}^{-7/6}\bigr)-1}\bigr]^{-1},$ (3)
$b\,$ $=\,\bigl[{\exp\bigl({0.51\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\bigl(1+0.69\sigma_{\mathrm{\scriptscriptstyle R}}^{2.4}\bigr)^{-5/6}}\bigr)-1}\bigr]^{-1},$ (4)
@@
where $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23\,C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ is the Rytov variance determining the atmospheric turbulence strength, $k_{\mathrm{w}}=2\pi/{\lambda_{\mathrm{w}}}$ is the optical wave number, $\lambda_{\mathrm{w}}$} is the wavelength, $L$ is the propagation distance, and $C_{\mathrm{n}}^{2}$ is the index of refraction structure parameter [14].
@@
Pointing Error: Pointing error refers to the misalignment of the laser beam with the optical receiver, caused by i) boresight, the fixed displacement between the beam centroid and the receiver center, and ii) jitter, the random temporal displacement due to factors such as building sway, wind, or vibrations. While careful deployment or fast-tracking transmitters can minimize boresight misalignment, residual jitter remains a concern even with a well-designed ATP subsystem [11]. To account for this effect, the severity of jitter in this work is later categorized into representative regimes reflecting different levels of ATP capability (see Table I in Section IV).
@@
After propagating a distance $L$, the laser beam with waist $w_{\mathrm{\scriptscriptstyle L}}$ is incident on the photodetector with aperture radius $r_{\mathrm{\scriptscriptstyle A}}$ [37]. The received optical beam attenuation due to turbulence-induced spread and pointing error, described by radial displacement $r$ between the beam footprint and the detector center, is given by $I_{\mathrm{p}}(r;L)\approx A_{\mathrm{\scriptscriptstyle 0}}\exp(-2r^{2}/w_{\mathrm{\scriptscriptstyle L_{\mathrm{eq}}}}^{2})$ where $w_{\mathrm{\scriptscriptstyle L_{\mathrm{eq}}}}^{2}=w_{\mathrm{\scriptscriptstyle L}}^{2}\sqrt{\pi}\,\text{erf}(v)/(2v\exp(-v^{2}))$ is the (squared) equivalent beam width with $v=\sqrt{\pi}r_{\mathrm{\scriptscriptstyle A}}/\sqrt{2}w_{\mathrm{\scriptscriptstyle L}}$. The fraction of the collected power at $r=0$ is given by
$A_{\mathrm{\scriptscriptstyle 0}}\,\approx\,\left[\mathrm{erf}(v)\right]^{2},$ (5)
@@
where $\text{erf}(\cdot)$ is the standard error function [36, Eq. (8.250.1)], and $0\leqslant A_{\mathrm{\scriptscriptstyle 0}}\leqslant 1$ [15]. For a coherent beam, the beam radius $w_{\mathrm{\scriptscriptstyle L}}$ is related to the beam waist $w_{\mathrm{\scriptscriptstyle 0}}$ at the optical transmitter’s exit aperture by $w_{\mathrm{\scriptscriptstyle L}}=w_{\mathrm{\scriptscriptstyle 0}}(1+\epsilon(\lambda_{\mathrm{w}}L/\pi w_{\mathrm{\scriptscriptstyle 0}}^{2})^{2})^{1/2}$ where $\epsilon=(1+2w_{\mathrm{\scriptscriptstyle 0}}^{2}/\rho_{\scriptscriptstyle 0}^{2}(L))$ is the global coherence parameter, $\rho_{\scriptscriptstyle 0}(L)=(1.46C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{2}L)^{-3/5}$ is the coherence length in turbulence [38], and the transmit beam waist $w_{\scriptscriptstyle 0}$ is related to the transmitter aperture diameter $D$ as $w_{\mathrm{\scriptscriptstyle 0}}={D}/{\sqrt{2}\pi}$ [39]. Farid and Hranilovic [15] derived the PDF of the irradiance component $I_{\mathrm{p}}$ under the assumptions of i.i.d. zero-mean Gaussian distributions for both horizontal and vertical receiver sways (i.e., zero boresight and identical jitters), expressed as
$\phantom{xxxxxxxxxx}f_{I_{\mathrm{p}}}(I_{\mathrm{p}})\,=\,\frac{\xi^{2}}{A_{\scriptscriptstyle 0}^{\xi^{2}}}I_{\mathrm{p}}^{\xi^{2}-1};\ \ \ 0\leqslant I_{\mathrm{p}}\leqslant A_{\scriptscriptstyle 0}\,$, (6)
@@
where the parameter $\xi\triangleq{w_{\mathrm{\scriptscriptstyle L_{eq}}}}/{2\sigma_{\mathrm{e}}}$ is the ratio of the received equivalent beam waist $w_{\mathrm{\scriptscriptstyle L_{eq}}}$ and the standard deviation of the pointing error displacement (jitter). This model is valid when $w_{\mathrm{\scriptscriptstyle L}}/r_{\mathrm{\scriptscriptstyle A}}\gg 1$ and further that the effects of $\xi$ and $A_{\scriptscriptstyle 0}$ parameters are independent [15].
@@
A recent refinement of the pointing error model is introduced in [40], where multiple alternative parameterizations of $\xi$ and $A_{\scriptscriptstyle 0}$ are proposed to enhance modeling accuracy and computational efficiency (see [40, Fig. 5]). In this work, we hence borrow the modified intensity uniform model from [40], which retains the same mathematical form for the probability distribution as in (6), but with redefined parameters given by
$A_{\mathrm{\scriptscriptstyle 0}}$ $\triangleq 1-\exp(-2r_{\mathrm{\scriptscriptstyle A}}^{2}/w_{\mathrm{\scriptscriptstyle L}}^{2}),$ (7)
$\xi^{2}$ $\triangleq r_{\mathrm{\scriptscriptstyle A}}^{2}/(2\sigma_{\mathrm{e}}^{2}A_{\mathrm{\scriptscriptstyle 0}}),$ (8)
Precisely, this choice demonstrates more than an order-of-magnitude reduction in normalized mean square error (NMSE) compared to the Farid–Hranilovic formulation (see [40, Table 1]), thereby offering a substantially more reliable characterization of pointing jitter effects in terrestrial FSO links.
@@
$f_{I}(I)=\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\,\Gamma(a)\Gamma(b)}\,G_{1,3}^{3,0}\left(\frac{abI}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,\,a-1,\,b-1\\ \end{array}\hskip-5.0pt\right).$ (9)
@@
II-B Noise Sources